What lump sum of money must be deposited into a bank account at the present time so that

​$400

per month can be withdrawn for

four

​years, with the first withdrawal scheduled for

five

years from​ today? The interest rate is

3​/4​%

per month.​ (Hint: Monthly withdrawals begin at the end of the month

60​.)
Answer #1

Monthly withdrawal = $400

No. of withdrawals = 48 ( for 4 years)

Withdrawal will start at the end of month 60.

R = 3/4% or .75%

So,

Present value of the withdrawals = (400*(1-1/1.0075^48)/.0075)*(1/1.0075^60)

Present value of the withdrawals = $10266.4 or $10266

Therefore, funds to be deposited in present value are $10266.4 or $10266.